OpenStudy (anonymous):

find the value of x \[\frac{ 1 }{ \frac{ 4 }{ 5 } } + \frac{ 1 }{ x } = \frac{ 1 }{ \frac{ 2 }{ 3 } }\]

OpenStudy (anonymous):

When dividing \[\frac{ 1 }{ \frac{ 1 }{ 2 } }\] the answer is 2, trying getting away from the complex fractions and then solve for x

OpenStudy (anonymous):

invert first\[\frac{ 5 }{ 4 }+x=\frac{ 3 }{ 2 }\]

OpenStudy (cwrw238):

no - don't invert the 1/x

OpenStudy (anonymous):

It doesn't work for the 2nd term \[\frac{ 1 }{ x }\neq x\]

OpenStudy (anonymous):

\[\frac{ 1 }{ \frac{ a }{ b } }=\frac{ b }{ a }\]

OpenStudy (anonymous):

oops sorry

OpenStudy (cwrw238):

lol - dont worry

OpenStudy (anonymous):

swept away with inverting

OpenStudy (anonymous):

It's fun to do! Looks alot nicer than 1/x, but we just can't do it...

OpenStudy (anonymous):

im atill stuck with what everybody has said lol

OpenStudy (anonymous):

If you divide by a fraction all you do is flip it upside down and multiply by it. So: \[\frac{1}{\frac{4}{5}}=1\times\frac{5}{4}=\frac{5}{4}\] Same with \(\frac{1}{\frac{2}{3}}=1\times\frac{3}{2}=\frac{3}{2}\) So using this we have: \[\frac{5}{4}+\frac{1}{x}=\frac{3}{2}\] Now multiply both sides of the equation by x to get: \[\frac{5}{4}x+1=\frac{3}{2}x\] Combine the \(x\) terms on the right hand side (subtract \(\frac{5}{4}x\) from both sides): \[1=\frac{1}{4}x\] Now multiply both sides by 4 to get: \[4=x\] Do you follow that, or can I explain any of it better?

OpenStudy (anonymous):

yes thats great thank you so much